Use the following cell phone airport data speeds​ (Mbps) from a particular network. Find

Upper P 40P40.

0.10.1

0.30.3

0.30.3

0.30.3

0.30.3

0.50.5

0.50.5

0.60.6

0.70.7

0.80.8

0.90.9

0.90.9

0.90.9

0.90.9

0.90.9

1.11.1

1.31.3

1.51.5

1.51.5

1.61.6

1.71.7

2.12.1

2.42.4

2.52.5

2.52.5

2.62.6

3.43.4

3.53.5

3.93.9

4.94.9

5.25.2

6.96.9

7.77.7

8.58.5

8.98.9

9.29.2

9.39.3

9.39.3

9.89.8

10.210.2

10.210.2

11.311.3

12.112.1

13.313.3

13.713.7

14.214.2

15.715.7

15.815.8

15.915.9

26.526.5

Upper P 40P40equals=nothing

Mbps
Answer #1

For P40, z- score = -0.2533

For the given data, μ = 5.582 and σ = 5.877

P40 = μ + z * σ = 5.582 - 0.2533 * 5.877 = 4.093